OpenStudy (anonymous):

circle with a diameter of 4, what is the circumference

OpenStudy (anonymous):

2*(pi)*4=8(pi)

OpenStudy (diyadiya):

Circumference = \(2 \times \pi \times radius \) \[radius = \frac{diameter}{2} \]

OpenStudy (anonymous):

srry 2*(pi)*2=4(pi)